# Optimal transport with a density-dependent cost function

### v1

@@ 1 Introduction
  • The map $T$ pairing two distributions is instrumental for the solution of matching problems [2];
  • The total cost $C$ provides a “horizontal” notion of distance between $p_{0,1}$, as opposed to “vertical” measures of discrepancy based on the pointwise comparison of their values, such as their $L^{2}$-distance and relative entropy [8];
  • When one of the distributions is known, the map $T$ estimates the other distribution through the change of variable formula. Similarly, when one of the distributions can be easily sampled, the map $T$ permits simulating the other, acting as a generative model [10];
  • Considering $y=T(x)$ as the end point of a path linking $x$ and $y$ allows one to interpolate between the two distributions [5];
  • The optimal transport problem is a critical building block for the more general optimal transport barycenter problem, which permits estimating and simulating conditional distributions [11].
@@
in normed spaces, may fail to capture some field-dependent notion of “true” distance between pairs of points. In Riemannian manifolds, for instance, a natural choice for $c$ is the squared geodesic distance between $x$ and $y$. Then the corresponding interpolation between distributions yields geodesics in Wasserstein space [7].
In problems in data analysis, it is natural to measure the distance between two points along the path joining them, provided that the areas that these paths traverse in phase space have a relatively large probability. It would not make sense to call two points “close” if connecting one to the other requires going through states that are unrealizable or highly unlikely. This was the logic used in [9] to define the Fermat distance between points, a data-based construction based on realizable paths.
@@
for which we impose the condition that the distributions underlying the $\{w(s^{j},0)\}$ and $\{x_{0}^{i}\}$ are the same, and similarly for $\{w(s^{j},1)\}$ and $\{x_{1}^{i}\}$. To enforce these, we penalize the difference between distributions through an instance of an adversarial procedure developed in [4] for the optimal transport barycenter problem.
@@ 1.1 Related work
Our choice of a cost function defined through an optimal path that avoids areas of low probability was inspired by the Fermat distance [9]. Yet the latter is heavily data-based –which is also one of its virtues– and hard to use as a cost for the optimal transport problem in a computational setting. A minimal action-based cost with density-dependent Lagrangian is much more suitable for this.
The literature on the OT problem starts with the classical work in Monge and Kantorovich [6, 3], separated by over a century, and continues afterward with very many significant contributions, which we cannot possibly summarize here. Most use explicitly defined cost functions, typically the squared Euclidean distance. Action-based costs appear naturally in the application of OT to Riemannian geometry (see [12, 13] and references therein), but we are not aware of work addressing their numerical implementation or making them depend on an ambient density. We hope that our work can help fill this gap and be of use in a geometrical context. Another connection between optimal transport and minimal action is the classical work in [1], which used a fluid mechanical interpretation to solve OT problems in a Lagrangian framework.
@@ 3 Optimal transport between two distributions
where the penalty function $R$ enforces alignment in distribution of the endpoint clouds with the empirical source/target sets. We determine $R$ through a methodology developed in [4], which we summarize here for the $\{w(s^{j},1)\}$ and $\{x_{1}^{i}\}$, since it applies without changes to the paths’ other end. We will assume for simplicity that $n_{0}=n_{1}=m$, i.e. that the number of samples available for the initial and final distributions and for the sample paths are the same (This is not required, since the samples must agree in distribution, not point-wise.)
@@ 5 Two applications
Applications of our procedure come in two main flavors. One uses the transportation cost $C(p_{0},p_{1})$ as a measure of dissimilarity between distributions. Such measures have wide applicability across data science; our illustration here uses them as a tool for grouping sets of distributions into clusters that are consistent with an ambient density $\rho_{0}$. The other flavor uses not the cost $C$ but the optimal map $y=T(x)$ itself; we illustrate this adopting $T$ as a tool for matching pairs of distributions, a task with applications in many fields, particularly in economics [2].
@@ 5.2 Matching points
Such order-preserving matching is exactly the pairing that our density-weighted action recovers. By contrast, the classical Wasserstein coupling with $c(x,y)=\tfrac{1}{2}\|x-y\|^{2}$ (i.e., Euclidean assignment) ignores the background density and thus uses chordal shortcuts across the disk, which produce a totally different pairing, “incorrect” from the perspective that considers the various distributions as rigid rotations of one another.
(a) New method (correct pairing).
(b) Wasserstein coupling (incorrect pairing).
Figure 5: Comparison of pairings: in this example, the proposed methodology recovers the intended matching between points, corresponding to a rigid rotation, while the Wasserstein endpoint coupling fails to pair them correctly.

### v2

@@ 1 Introduction
  • The map $T$ pairing two distributions is instrumental for the solution of matching problems [Galichon2016];
  • The total cost $C$ provides a “horizontal” notion of distance between $p_{0,1}$, as opposed to “vertical” measures of discrepancy based on the pointwise comparison of their values, such as their $L^{2}$-distance and relative entropy [Santambrogio2015];
  • When one of the distributions is known, the map $T$ estimates the other distribution through the change of variable formula. Similarly, when one of the distributions can be easily sampled, the map $T$ permits simulating the other, acting as a generative model [TabakTrigila2016];
  • Considering $y=T(x)$ as the end point of a path linking $x$ and $y$ allows one to interpolate between the two distributions [McCann1997];
  • The optimal transport problem is a critical building block for the more general optimal transport barycenter problem, which permits estimating and simulating conditional distributions [TabakTrigila2018].
@@
in normed spaces, may fail to capture some field-dependent notion of “true” distance between pairs of points. In Riemannian manifolds, for instance, a natural choice for $c$ is the squared geodesic distance between $x$ and $y$. Then the corresponding interpolation between distributions yields geodesics in Wasserstein space [Otto2001].
In problems in data analysis, it is natural to measure the distance between two points along the path joining them, provided that the areas that these paths traverse in phase space have a relatively large probability. It would not make sense to call two points “close” if connecting one to the other requires going through states that are unrealizable or highly unlikely. This was the logic used in [SapienzaGroismanJonckheere2018] to define the Fermat distance between points, a data-based construction based on realizable paths.
@@
for which we impose the condition that the distributions underlying the $\{w(s^{j},0)\}$ and $\{x_{0}^{i}\}$ are the same, and similarly for $\{w(s^{j},1)\}$ and $\{x_{1}^{i}\}$. To enforce these, we penalize the difference between distributions through an instance of an adversarial procedure developed in [LipnickTabakTrigilaWangYeZhao2025] for the optimal transport barycenter problem.
@@ 1.1 Related work
Our choice of a cost function defined through an optimal path that avoids areas of low probability was inspired by the Fermat distance [SapienzaGroismanJonckheere2018]. Yet the latter is heavily data-based –which is also one of its virtues– and hard to use as a cost for the optimal transport problem in a computational setting. A minimal action-based cost with density-dependent Lagrangian is much more suitable for this.
The literature on the OT problem starts with the classical work in Monge and Kantorovich [Monge1781, Kantorovich1942], separated by over a century, and continues afterward with very many significant contributions, which we cannot possibly summarize here. Most use explicitly defined cost functions, typically the squared Euclidean distance. Action-based costs appear naturally in the application of OT to Riemannian geometry (see [Villani2009, Villani2003] and references therein), but we are not aware of work addressing their numerical implementation or making them depend on an ambient density. We hope that our work can help fill this gap and be of use in a geometrical context. Another connection between optimal transport and minimal action is the classical work in [BenamouBrenier2000], which used a fluid mechanical interpretation to solve OT problems in a Lagrangian framework.
@@ 3 Optimal transport between two distributions
where the penalty function $R$ enforces alignment in distribution of the endpoint clouds with the empirical source/target sets. We determine $R$ through a methodology developed in [LipnickTabakTrigilaWangYeZhao2025], which we summarize here for the $\{w(s^{j},1)\}$ and $\{x_{1}^{i}\}$, since it applies without changes to the paths’ other end. We will assume for simplicity that $n_{0}=n_{1}=m$, i.e. that the number of samples available for the initial and final distributions and for the sample paths are the same (This is not required, since the samples must agree in distribution, not point-wise.)
@@ 5 Two applications
Applications of our procedure come in two main flavors. One uses the transportation cost $C(p_{0},p_{1})$ as a measure of dissimilarity between distributions. Such measures have wide applicability across data science; our illustration here uses them as a tool for grouping sets of distributions into clusters that are consistent with an ambient density $\rho_{0}$. The other flavor uses not the cost $C$ but the optimal map $y=T(x)$ itself; we illustrate this adopting $T$ as a tool for matching pairs of distributions, a task with applications in many fields, particularly in economics [Galichon2016].
@@ 5.2 Matching points
We can see in figure (5) the results of applying both our density-weighted action methodology and classical Wasserstein coupling with $c(x,y)=\tfrac{1}{2}\|x-y\|^{2}$ (i.e., Euclidean assignment). As we can see, the former gives results much closer to the order-preserving matching that the intuition above suggests than the latter, which produce a totally different pairing, ignoring the background density and thus using chordal shortcuts across the disk. The pairing provided by the new procedure does not agree exactly with the “ground truth” suggested by intuition, since the paths of minimal action are not isopycnal (i.e. lines of constant density), yet it approximates it to a high degree.
(a) New method.
(b) Wasserstein coupling.
Figure 5: Comparison of pairings: in this example, the proposed methodology approximates the intended matching between points, corresponding to a rigid rotation, while the Wasserstein endpoint coupling does that intuition.